Riverdale Hospital is concerned over the number of times patients must wait more than 30 minutes beyond their scheduled appointments. The quality team decided to take random samples of 75 patients to see how many in each sample had to wait more than 30 minutes. Each instance is considered a defect in the clinic process. The table below contains the data for 15 samples.

Sample Number of defects
1 5
2 2
3 1
4 3
5 1
6 5
7 2
8 3
9 6
10 3
11 9
12 9
13 5
14 2
15 3

LaTeX: \sigmaσp = ______ (enter response rounded to 3 decimal places).
Answer #1

Solution:

\bar{p}= No. of Defectives/ Total Inspected

No. of Defectives=59

Total Inspected=75*15=1125

So \bar{p}= 59/1125= 0.052

P(1 - 1) Op = V!

here n= Sample Size= 75

p(1-7) Op = V n 0.052(1 -0.052) 75

\sigma _{p} =0.026